OpenStudy (sonyalee77):

Laura is completing a craft project that involves covering only the lateral surface of a cylindrical container with fabric. The cylinder has a height of 9.7 in. and a radius of 10.8 in. To the nearest square unit, how much fabric does she need for this project? Use a calculator.

OpenStudy (rajee_sam):

Lateral surface is the Area of sides only. The top and the bottom are not included. So the Total Surface Area of a Cylinder is given by the formula: \[(2\times \pi \times r ^{2}) +( 2 \times \pi \times r \times h)\] now the \[2 \times \pi \times r ^{2}\] is the Area of the top and the bottom. So Lateral Surface area is just \[2 \times \pi \times r \times h\] Just plug in the values for r and h given to you and find your answer.